Find equation to the tangent at A to the circle (x-a)^2 + (y-b)^2 =r^2 where the radius through A makes an angle alpha with the x axis.

Grade:12

1 Answers

Ajay
7 years ago
The Co-ordinates of center are (a,b) and radius is r.
Since radius to A makes angle \alpha with x axis the co-ordinates of A are
(a + rcos\alpha, b+rsin\alpha).
Now we know A its easy to find equation of tangent at A